# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

### v1

@@ 3.5. Participants
We recruited 600 participants from Prolific, divided equally among the four prediction displays, ensuring a gender-balanced sample. We set additional screening criteria so that all of our participants are (1) based in the USA; (2) no vision impairment; (3) no colorblindness; (4) English as the first language; and (5) aged between 18 and 65. This led to 150 participants in each condition representing different prediction displays. We determined the target sample size using a nonparametric simulation of the experiment bootstrapped from a pilot study of 30 participants. By grouping trial-level responses according to all manipulated factors and bootstrapping these responses with replacement, we simulated trial-level outcomes. We identified the target sample size (600) that resulted in confidence intervals on accuracy with width less than .
@@ 3.6. Analysis Method
We pre-registered an analysis plan focusing on three response variables: (1) accuracy, (2) shortest path length, and (3) willingness-to-pay, and followed our pre-registered plan in all analyses reported below unless otherwise specified. We define accuracy (ACC) as binary, indicating whether a participant selected the correct label for the task. Shortest path (SP) length is a continuous variable representing the number of hops from the leaf node a participant selects to the leaf node of the ground truth label. SP length is a measure widely used to quantify the semantic similarity (Resnik, 1995; Budanitsky and Hirst, 2006) between categories in a taxonomy. Although highly correlated with accuracy, this measure provides further information on how “incorrectness” varied by our experimental manipulations. Willingness-to-pay (WTP) ranges from and , and measures how much they will pay to have access to the display.
@@
We pre-registered Bayesian Linear Mixed-effect models for accuracy and shortest path length. Our model specifications are presented in Model 1 and Model 2. In this context, trial is numeric, indicating the trial number, whereas condition (4 levels: baseline, Top-, Top-, and prediction set), shift (2 levels: in-distribution and OOD), difficulty (2 levels: easy and hard), and size (2 levels: small and large) are factors corresponding to our experimental manipulations. The term ID represents each participant’s unique Prolific ID.
@@
For the accuracy model, we use a Bernoulli distribution for the likelihood shown in Line 1 left, parameterized by , which denotes the probability of a correct label. Line 3 presents the hierarchical logistic regression model. In this specification, we model the logit of through interactions among all fixed effects and incorporate random intercepts and slopes for trial, grouped by participants’ unique ID. This approach allows us to account for participants’ baseline performance differences and variations in rates of change across conditions and trials.
@@
For the shortest path length model, we employ a Zero-inflated Negative Binomial (ZINB) model as the likelihood, parameterized by (expected SP length), (negative binomial dispersion), and (zero-inflation probability), as shown in Line 1 right. The expected path length between the participants’ responses and the truth label is modeled by the negative binomial component in Line 3. We apply a log-link function to , modeled by all fixed effects along with their interactions with random intercepts and slopes for trials grouped by participant ID. We do not explicitly model the dispersion parameter for the NB component. However, we placed an exponential prior with a rate of to moderate the degree of dispersion in our model. Given that our observed data had a maximum SP length of 27, this prior acts as a regularizing component, ensuring our model captures the dispersion of the observed data without predicting implausibly large SP lengths. Line 5 shows the zero-inflation component that models the proportion of zero path length (i.e., correct answer) not captured by the NB components. We apply a logit link function to and model it by all fixed effects along with their interactions.
@@
We deviated from our pre-registered model specifications only in setting the priors. We originally aimed to set weakly informative priors, but several were too narrow to allow model convergence. We therefore widened the priors to be even less informative. Full details are available in supplemental material.
@@
We followed a standard Bayesian workflow (Gelman et al., 2020) to check if both models fit. Detailed model diagnostics are included in the Supplemental Material. We use the models to generate predictions of the respective response variable based on 1,000 draws from the posterior distribution of the models’ fixed effects for each unique combination of our experimental manipulations while holding the trial effect constant at the mean. Our results present the expected predictions as point estimates, with uncertainty expressed through highest posterior density interval (HDI).
@@
Specifically, we first calculate the expected bonus participants gained in each condition by multiplying the expected accuracy rates by the maximum earnable bonus of . We then calculate the expected bonus difference with access to each prediction display relative to the baseline by subtracting the expected bonus of the baseline from the expected bonus of each treatment conditions. We finally evaluate the discrepancy between this expected bonus difference and the corresponding WTPs reported by participants who used each prediction display, by constructing a “willingness-to-overpay” metric. This measure reflects the difference between the additional bonus one is expected to earn from having access to a prediction display over the baseline and the actual difference in the average bonus participants received from accessing a prediction display relative to the baseline. We use bootstrapping to derive 95% confidence intervals with the median as a point estimate.
@@ 3.7. Qualitative Analysis of Strategies
We performed a qualitative analysis and developed a bottom-up open coding scheme using the strategies provided by the participants. Two coders worked in collaboration, with each coding half of the strategies. After coding, any ambiguous strategies were discussed until a mutual agreement was reached. Among 585 valid strategies, we excluded 41 uninformative ones (Code: 1) with responses such as “I have no idea, trying my hardest”. For the remaining strategies, we identified 17 ways in which the prediction displays and the search features were used. They are summarized by the categories below.
How do participants use different prediction displays? (Code: 2–6)
Do participants find predictions to be trustworthy? (Code: 7–10)
How do participants use the search feature to find a preferred label? (Code: 11–15)
How do participants narrow down their choices and identify a preferred label? (Code: 16–18)
@@ 4.1. Data Preliminaries
We received 599 valid responses after excluding one participant from the baseline condition according to our pre-registered exclusion rule. As demonstrated in Figure 5 left, on average, participants spent minutes completing all 16 image labeling tasks for all conditions. Participants in Top-10 and RAPS conditions took slightly longer to complete than those of the baseline and Top-1 conditions.
Distribution of participants’ study completion time and the amount of bonus earned.
From the distribution of earned bonus shown in Figure 5 right, we see a clear difference that having access to prediction displays can help participants get more correct answers and hence a higher bonus reward than those who did not have access to the predictions. On average, baseline participants provided seven correct labels, which is lower than those who used prediction displays (10 out of 16 correct).
@@ 4.2.1. Overview
In Figure 6, we present the expected predictions by our accuracy model (Model 1) for each type of image stimuli, with uncertainty quantified as HDIs. At a high level, we find that RAPS does not improve participants’ labeling accuracy for easy tasks, but is more useful for hard tasks, especially for OOD images. While a smaller RAPS can be more useful for in-distribution images, size does not appear to affect accuracy much when images are OOD. Because small and large set sizes depend on whether instances are in-distribution or OOD in our data-generating model (Section 3.2), we provide the mean set size for size groups to contextualize results below.
Accuracy results
@@ 4.2.2. In-distribution
This effect is most prominent when images are classified as hard in Figure 6, (C) and (D). When using RAPS to label hard images, participants’ accuracy is higher when set size is smaller (; HDI: ; mean size: ), and lower when set size is larger (; HDI: ; mean size: ). We speculate the reason is that a larger set size increases participants’ cognitive load as they need to traverse through the set and evaluate all labels. Figure 6 (C) shows that expected accuracy of RAPS participants is roughly % higher than that of the Top-10 condition (; HDI: ) which is roughly % higher than the Top-1 condition (; HDI: ), though HDIs overlap.
@@
This is particularly evident for in-distribution images categorized as easy (cf. Figure 6, (A) and (B)). Here, the participants’ labeling accuracy tends to be lower than the prediction display accuracy (i.e., the percentage of the time that the prediction display included the true label), depicted as solid horizontal lines. However, when labeling hard images, participants in the Top-1 condition can improve upon the model’s prediction. As suggested by our open codes discussed in Section 4.5, we find when the Top-1 prediction was obviously inaccurate, participants were more likely to rely on their own judgment and used the search tool to identify the correct answer. The presence of a greater number of labels in the Top-10 and RAPS conditions may have complicated the labeling process due to increased cognitive load. This is exemplified by the observed trends in these conditions: among incorrect responses, participants in the RAPS condition missed the correct label from the prediction set of the time, whereas those in the Top-10 condition did so of the time.
In summary, when task images are in-distribution, (1) labeling accuracy for participants in the Top-1 or Top-10 conditions was higher when the images are easy (cf. Figure 6 (A) and (B)); (2) the accuracy of RAPS participants is negatively correlated with set size: a larger set size decreases accuracy, particularly for hard images (cf. Figure 6 (C) and (D)); and (3) Top-1 participants are more likely to rely on their own judgment and use the search tool to identify the correct answer when predictions were less likely to be correct. In contrast, the pattern of incorrect responses observed in the Top-10 and RAPS conditions suggests that the presence of more predictions in the display made it harder for participants to discern between inaccurate and accurate predictions, even when the correct label was present.
@@ 4.2.3. Out-of-distribution
Due to the similarity in HDIs across conditions for easy OOD images (cf. Figure 6 (E) and (F)), we report expected accuracy marginalized over set sizes. We find Top-1 participants tend to achieve slightly higher accuracy in expectation (; HPI: ) than those using Top-10 (; HDP: ). Meanwhile, the RAPS condition yielded the lowest expected accuracy of (HPI: ), despite an average set size of (as detailed in Table 3) that is smaller than the size of Top-10, with identical chance that the prediction display contained the true label (colored lines in Figure 6 (E) and (F)). This divergence from our in-distribution finding, where smaller set sizes typically correlate with higher labeling accuracy when predictions are reliable, suggests additional factors may be at play in the OOD scenario. It is possible the reduction in cognitive load afforded by access to softmax pseudo-probabilities in the Top-10 condition for use in prioritizing labels aided participants in finding the correct label. This observation partially reinforces our finding from the easy in-distribution results: when the model is well-calibrated (cf. Figure 6 (E) and (F)), viewing fewer predictions and having access to prediction confidence (i.e., as in our Top- conditions) can improve accuracy.
When OOD images are hard, RAPS yield the highest labeling accuracy regardless of set size
For hard OOD images, RAPS set sizes grow substantially larger, with small sets averaging instances and large sets averaging instances. When hard OOD images had smaller set size (Figure 6 (G)), RAPS participants outperformed those in Top- conditions with a relatively high accuracy of (HPI: ), followed by the Top-10 condition (; HPI: ) and Top-1 condition (; HPI: ). When set size was larger (Figure 6 (H)), we identified a consistent pattern that RAPS participants (, HPI: ) consistently outperformed those in the Top- conditions. Nevertheless, the accuracy of RAPS participants is significantly lower than the average coverage rate (Figure 6 (G) and (F) green solid horizontal lines), implying that even when the true label is included in the set, it is still challenging for participants to identify it.
@@
For hard OOD images with a smaller average set size (Figure 6, (G)), we find that participants in the baseline condition who did not have access to any predictions have a higher accuracy of (HPI: ) than those in the Top-1 condition (; HPI: ). For hard OOD images with a larger average set size (Figure 6, (H)), the accuracy of using RAPS (, HPI: ) versus the baseline condition (; HPI: ) is similar, and participants who used Top-1 prediction are more accurate (; HPI: ) than those who used Top-10 predictions (; HPI: ). Our results suggest that Top- predictions from a poorly calibrated model (cf. Figure 6 (G) and (H)) can negatively influence human judgment. Although RAPS participants had relatively higher accuracy than those in the Top- conditions, their accuracy is almost identical to those in baseline, suggesting it might be beneficial to withhold predictions for hard OOD stimuli that require a large set size to achieve the desired coverage.
In summary, (1) when OOD images are easy and the predictions and softmax scores well-calibrated (cf. Figure 6 (E) and (F)), Top- predictions tend to result in higher labeling accuracy; (2) when OOD images are hard and the model predictions more error-prone (cf. Figure 6 (G) and (H)), RAPS participants can outperform their Top- counterparts. Unlike in the in-distribution case, increasing the size of the prediction set does not degrade accuracy; and (3) when OOD task images are hard and the predictions more error-prone, participants who rely on their own judgment and the label space search tool may be more likely to identify the correct answer. The reasons behind the performance advantage of participants in the baseline condition when labeling hard OOD images with larger set sizes remain ambiguous. It is possible the accuracy decrease for participants in Top-10 condition when labeling hard OOD images that have larger set size is due to overreliance on inaccurate predictions, given that the average coverage for Top-10 predictions decreases from (small) to (large), as shown in Table 3.
@@ 4.3. Shortest Path Length
The predictions of our SP length model (Model 2) highly correlate with those of the accuracy model when comparing the results in Figure 6 with those in Figure 7. We briefly highlight the main takeaways that complement our accuracy results, with more detailed descriptions in Appendix B.1.
Shortest path length results
@@
Recall that in Section 4.2.3, we find when OOD images are classified as hard, RAPS participants can outperform their Top- counterparts, but there is no difference in accuracy by set size (cf. Figure 6, (G) and (H)). However, from the SP length, we find if the task image has a larger set size, RAPS participants’ chosen label is slightly closer to the correct answer in the label space hierarchy. As demonstrated in Figure 7 (G) and (H), the SP length for hard OOD images that have a larger set is (HPI: , mean size: ) and that of a smaller set is 4.2 (HPI: ; mean size: ). It is difficult to say why this might be the case; it is possible, as supported by our open codes in Section 4.5, that when OOD stimuli are especially challenging to recognize, participants spend more time scrutinizing labels in the prediction set clues, and that even when the probability that the prediction set contains the true label is the same, viewing more predictions helps cue their ability to identify the true label. The effect is quite small, whatever the cause.
@@
In summary, we find (1) for hard OOD stimuli, although the labeling accuracy of RAPS participants is similar, a larger prediction set may guide responses closer to the ground truth (cf. Figure 7 (G) and (H)); and (2) when the model is well calibrated, inaccurate Top- predictions might still guide participants to identify the correct label.

### v6

@@ 3.5. Participants
We recruited 600 participants from Prolific, divided equally among the four prediction displays, ensuring a gender-balanced sample. We set additional screening criteria so that all of our participants were (1) based in the USA, (2) had no vision impairment, (3) had no colorblindness, (4) spoke English as the first language, and (5) aged between 18 and 65. This led to 150 participants in each condition representing different prediction displays. We determined the target sample size using a non-parametric simulation of the experiment bootstrapped from a pilot study of 30 participants. By grouping trial-level responses according to all manipulated factors and bootstrapping these responses with replacement, we simulated trial-level outcomes. We identified the target sample size (600) that resulted in confidence intervals on accuracy with widths less than .
@@ 3.6. Analysis Method
We pre-registered an analysis plan focusing on three response variables: (1) accuracy, (2) shortest path length, and (3) willingness-to-pay, and followed our pre-registered plan in all analyses reported below unless otherwise specified. We define accuracy (ACC) as binary, indicating whether a participant selected the correct label for the task. The shortest path (SP) length is a continuous variable representing the number of hops from the leaf node a participant selects to the leaf node of the ground truth label. SP length is a measure widely used to quantify semantic similarity (Resnik, 1995; Budanitsky and Hirst, 2006) between categories in a taxonomy. Although highly correlated with accuracy, this measure provides further information on how “incorrectness” varied by our experimental manipulations. Willingness-to-pay (WTP) ranges from to , measuring how much they will pay to access the display.
@@
We pre-registered Bayesian Linear Mixed-effect models for accuracy and SP length. Our model specifications are presented in Model 1 and Model 2. In this context, trial is numeric, indicating the trial number, whereas condition (4 levels: baseline, Top-, Top-, and prediction set), shift (2 levels: in-distribution and OOD), difficulty (2 levels: easy and hard), and size (2 levels: smaller and larger) are factors corresponding to our experimental manipulations. The term ID represents each participant’s unique Prolific ID.
@@
For the accuracy model stated in Model 1, we use a Bernoulli distribution for the likelihood shown in Line 1, parameterized by , which denotes the probability of a correct label. Line 1 presents the hierarchical logistic regression model. In this specification, we model the logit of through interactions among all fixed effects and incorporate random intercepts and slopes for trial, grouped by participants’ unique ID. This approach allows us to account for participants’ baseline performance differences and variations in rates of change across conditions and trials.
@@
For the SP length model stated in Model 2, we employ a Zero-inflated Negative Binomial (ZINB) model as the likelihood, parameterized by (expected SP length), (negative binomial dispersion), and (zero-inflation probability), as shown in Line 2. The expected path length between the participants’ responses and the truth label is modeled by the negative binomial component in Line 2. We apply a log-link function to , modeled by all fixed effects along with their interactions with random intercepts and slopes for trials grouped by participant ID. We do not explicitly model the dispersion parameter for the NB component. However, we placed an exponential prior with a rate of to moderate the degree of dispersion in our model. Given that our observed data had a maximum SP length of 27, this prior acts as a regularizing component, ensuring our model captures the dispersion of the observed data without predicting implausibly large SP lengths. Line 2 shows the zero-inflation component that models the proportion of zero path length (i.e., correct answer) not captured by the NB components. We apply a logit link function to and model it by all fixed effects along with their interactions.
@@
We deviated from our pre-registered model specifications only in setting the priors. We initially aimed to set weakly informative priors, but several were too narrow to allow model convergence. We therefore widened the priors to be even less informative. Full details are available in the Supplemental Material.
@@
We followed a standard Bayesian workflow (Gelman et al., 2020) to check if both models fit. Detailed model diagnostics are included in the Supplemental Material. We use the models to generate predictions of the respective response variable based on 1,000 draws from the posterior distribution of the models’ fixed effects for each unique combination of our experimental manipulations while holding the trial effect constant at the mean. Our results present the expected predictions as point estimates, with uncertainty expressed through the highest posterior density interval (HDI).
@@
Specifically, we first calculate the expected bonus participants gained in each condition by multiplying the expected accuracy rates by the maximum earnable bonus of . We then calculate the expected bonus difference with access to each prediction display relative to the baseline by subtracting the expected bonus of the baseline from the expected bonus of each treatment condition. We finally evaluate the discrepancy between this expected bonus difference and the corresponding WTPs reported by participants who used each prediction display by constructing a “willingness-to-overpay” metric. This measure reflects the difference between the additional bonus one is expected to earn from having access to a prediction display over the baseline and the actual difference in the average bonus participants received from accessing a prediction display relative to the baseline. We use bootstrapping to derive 95% confidence intervals with the median as a point estimate.
@@ 3.7. Qualitative Analysis of Strategies
We performed a qualitative analysis and developed a bottom-up open coding scheme using the strategies provided by the participants. Two coders worked in collaboration, each coding half of the strategies. After coding, any ambiguous strategies were discussed until mutual agreements were reached. Among 585 valid strategies, we excluded 41 uninformative ones (Code: 1) with responses such as “I have no idea, trying my hardest”. For the remaining strategies, we identified 17 ways in which the prediction displays and the search features were used, as illustrated in Table 4.
| Code | Category |
| 2–6 | How do participants use different prediction displays? |
| 7–10 | Do participants find predictions to be trustworthy? |
| 11–15 | How do participants use the search feature to find a preferred label? |
| 16–18 | How do participants narrow down their choices and identify a preferred label? |
@@ 4.1. Data Preliminaries
We received 599 valid responses after excluding one participant from the baseline condition according to our pre-registered exclusion rule. As demonstrated in Figure 5 left, participants spent, on average, minutes completing the 16 image labeling tasks. Participants in the Top-10 and RAPS conditions took slightly longer to complete tasks than those in the baseline and Top-1 conditions.
This figure presents the distributions of participants’ study completion time and the amount of bonus earned by treatment conditions. Each distribution has a solid vertical line showing the average and two dotted vertical lines showing the interquartile range.
From the distribution of earned bonuses shown in Figure 5 right, there is a clear difference indicating that access to the prediction displays helped participants achieve more correct answers and hence a higher bonus reward than those without access. On average, baseline participants provided seven correct labels, which is lower than those who used prediction displays (10 out of 16 correct).
@@ 4.2.1. Overview
In Figure 6, we present the expected predictions by our accuracy model (Model 1) for each type of image stimuli, with uncertainty quantified as HDIs. At a high level, we find that RAPS does not improve participants’ labeling accuracy for easy tasks but is more useful for hard tasks, especially for OOD images. While a smaller RAPS can be more useful for in-distribution images, size does not appear to affect accuracy much when images are OOD. Because smaller and larger set sizes depend on whether instances are in-distribution or OOD in our data-generating model (Section 3.2), we provide the average set size for size groups to contextualize the results below.
This figure displays a grid presenting the median expected labeling accuracy predicted by our accuracy model with uncertainty expressed as 95% HDIs. Rows differentiate in-distribution and OOD results, while columns separate results by difficulty and set size. Solid horizontal lines for each stimuli group illustrate the prediction accuracy achieved by that display (i.e., the probability that the displayed predictions included the true label) using our base classifier (i.e., WRN).
@@ 4.2.2. In-distribution
This effect is most prominent when images are classified as hard in Figure 6, (C) and (D). When using RAPS to label hard images, participants’ accuracy is higher when the set size is smaller (; HDI: ; average size: ), and lower when set size is larger (; HDI: ; average size: ). We speculate that the reason is that a larger set size increases participants’ cognitive load as they need to traverse through the set and evaluate all labels. Figure 6 (C) shows that expected accuracy of RAPS participants is roughly % higher than that of the Top-10 condition (; HDI: ) which is roughly % higher than the Top-1 condition (; HDI: ), though HDIs overlap.
@@
This is particularly evident for in-distribution images categorized as easy (cf. Figure 6, A and B). Here, the participants’ labeling accuracy tends to be lower than the prediction display accuracy (i.e., the percentage of the time the prediction display included the true label), depicted as solid horizontal lines. However, when labeling hard images, participants in the Top-1 condition can improve upon the model’s prediction. As suggested by our open codes discussed in Section 4.5, we find that when the Top-1 prediction was obviously inaccurate, participants were more likely to rely on their own judgment and used the search tool to identify the correct answer. The presence of a greater number of labels in the Top-10 and RAPS conditions may have complicated the labeling process due to an increased cognitive load. This is exemplified by the observed trends in these conditions: among incorrect responses, participants in the RAPS condition missed the correct label from the prediction set of the time, whereas those in the Top-10 condition did so of the time.
In summary, when task images are in-distribution, (1) labeling accuracy for participants in the Top-1 or Top-10 conditions is higher when the images are easy (cf. Figure 6, A and B); (2) the accuracy of RAPS participants is negatively correlated with set size: a larger set size decreases accuracy, particularly for hard images (cf. Figure 6, C and D); and (3) Top-1 participants are more likely to rely on their own judgment and use the search tool to identify the correct answer when predictions are less likely to be correct. In contrast, the pattern of incorrect responses observed in the Top-10 and RAPS conditions suggest that the presence of more predictions makes it harder for participants to discern between inaccurate and accurate predictions, even when the correct label is present.
@@ 4.2.3. Out-of-distribution
Due to the similarity in HDIs across conditions for easy OOD images (cf. Figure 6, E and F), we report expected accuracy marginalized over set sizes. We find Top-1 participants tend to achieve slightly higher accuracy in expectation (; HPI: ) than those using Top-10 (; HDP: ). Meanwhile, the RAPS condition yields the lowest expected accuracy of (HPI: ), despite an average set size of (as detailed in Table 3) that is smaller than the size of Top-10, with identical chance that the prediction display contained the true label (i.e., colored horizontal lines in Figure 6, E and F). This divergence from our in-distribution finding, where smaller set sizes typically correlate with higher accuracy when predictions are reliable, suggests that additional factors may be at play in the OOD scenario. It is possible that the reduction in cognitive load afforded by having access to softmax scores in the Top-10 condition aided participants in finding the correct label. This observation partially reinforces our finding from the easy in-distribution results: when the model is well-calibrated (cf. Figure 6, E and F), viewing fewer predictions and having access to prediction confidence (i.e., as in our Top- conditions) can improve accuracy.
This figure displays a grid presenting the median expected shortest path length predicted by our shortest path model with uncertainty expressed as 95% HDIs. Rows differentiate in-distribution and OOD results, while columns separate results by difficulty and set size.
When OOD images are hard, RAPS yield the highest labeling accuracy regardless of the set size
For hard OOD images, RAPS set sizes grow substantially larger, with smaller sets averaging instances and larger sets averaging instances. When hard OOD images had smaller set sizes (i.e., Figure 6, G), RAPS participants outperform those in Top- conditions with a relatively high accuracy of (HPI: ), followed by the Top-10 condition (; HPI: ) and Top-1 condition (; HPI: ). When the set size is larger (i.e., Figure 6, H), we identify a consistent pattern that RAPS participants (, HPI: ) outperform those in the Top- conditions. Nevertheless, the accuracy of RAPS participants is still lower than the average coverage rate (cf. Figure 6, G and F, with green solid horizontal lines), implying that even when the sets include the true label, it is still challenging for participants to identify it.
@@
For hard OOD images with smaller average set size (i.e., Figure 6, G), we find that participants in the baseline condition who did not have access to any predictions have a higher accuracy of (HPI: ) than those in the Top-1 condition (; HPI: ). For hard OOD images with a larger average set size (i.e., Figure 6, H), the accuracy of using RAPS (, HPI: ) versus the baseline condition (; HPI: ) is similar, and participants who used Top-1 prediction are more accurate (; HPI: ) than those who used Top-10 predictions (; HPI: ). Our results suggest that Top- predictions from a poorly calibrated model (cf. Figure 6, G and H) can negatively influence human judgment relative to no model assistance. Although RAPS participants have relatively higher accuracy than those in the Top- conditions, their accuracy is almost identical to those in baseline, suggesting it might be beneficial to withhold predictions for hard OOD stimuli that require a larger set size to achieve the desired coverage, at least in an OOD setting like ours, where the perturbations decreased both human and AI performance.
In summary, (1) when OOD images are easy and the predictions and softmax scores well-calibrated (cf. Figure 6, E and F), Top- predictions tend to result in higher labeling accuracy; (2) when OOD images are hard and the model predictions more error-prone (cf. Figure 6, G and H), RAPS participants can outperform their Top- counterparts. Unlike in the in-distribution case, increasing the size of the prediction set does not degrade accuracy; and (3) when OOD images are hard and the model predictions are more error-prone, participants who rely on their own judgment and the label space search tool may be more likely to identify the correct answer. The reasons behind the performance advantage of the participants in the baseline condition when labeling hard OOD images with larger set sizes remain ambiguous. It is possible that the accuracy decrease for participants in Top-10 condition when labeling hard OOD images that have larger set size is due to overreliance on inaccurate predictions, given that the average coverage for Top-10 predictions decreases from (smaller) to (larger), as shown in Table 3.
@@ 4.3. Shortest Path Length
The predictions of our SP length model (Model 2) are highly correlated with those of the accuracy model when comparing the results in Figure 6 with those in Figure 7. We briefly highlight the main takeaways that complement our accuracy results, with more detailed descriptions in Appendix B.2.
@@
Recall that in Section 4.2.3, we find when OOD images are classified as hard, RAPS participants can outperform their Top- counterparts, but there is no difference in accuracy by set size (cf. Figure 6, G and H). However, from the SP length, we find that if the task image has a larger set size, RAPS participants’ chosen label is slightly closer to the correct answer in the label space hierarchy. As demonstrated in Figure 7 (G) and (H), the SP length for hard OOD images that have a larger set size is (HPI: , average size: ) and that of a smaller set size is 4.2 (HPI: ; average size: ). It is difficult to say why this might be the case; it is possible, as supported by our open codes in Section 4.5, that when OOD stimuli are especially challenging to recognize, participants spend more time scrutinizing labels in the prediction set for clues, and that even when the probability that the prediction set contains the true label is the same, viewing more predictions helps cue their ability to identify the true label. The effect is quite small, whatever the cause.
@@
In summary, we find (1) for hard OOD stimuli in our setting, although the accuracy of RAPS participants is similar, a larger set may guide responses closer to the ground truth (cf. Figure 7, G and H); and (2) when the model is well-calibrated, inaccurate Top- predictions might help participants in identifying the correct label.